OpenStudy (anonymous):

Solve (1/x) = (6/3x) + 1

OpenStudy (ivettef365):

you can write the equation as follows: 1 6 ---- = === + 1 x 3x

OpenStudy (ivettef365):

then you can multiply both sides by "x" and you have x . 1 6 --- = ---- +1 (x) x 3x

OpenStudy (ivettef365):

do you know how to solve this

OpenStudy (anonymous):

no

OpenStudy (ivettef365):

ok, on the left side the x . 1/x eliminates the x so you have and I will multiply the x on the right hand side and you get 1 = 6x + x --- 3x

OpenStudy (ivettef365):

please let me know if you understand what I did so far

OpenStudy (dumbcow):

6/3 = 2 so you can say \[\frac{1}{x} = \frac{2}{x} +1\] write "1" as a fraction \[\frac{1}{x} = \frac{2}{x} +\frac{x}{x}\] then \[1 = 2+x\] ...